OpenStudy (anonymous):

Using complete sentences, explain how to completely factor 2x^2 − 8x − 42.

OpenStudy (anonymous):

First, take a two out of the equation to make it simpler. The equation then becomes x^2 - 4x - 21. Next, find two numbers that both multiply to make -21 and add up to make -4. The numbers -7 and 3 do exactly that. Next - factor them by placing the numbers with an x in parentheses (x+3) (x-7). Lastly, check your work by foiling!

jhonyy9 (jhonyy9):

yes is right but you have missed the firstly 2 and hence with this 2 your equation will be factored 2(x+3)(x-7)

jhonyy9 (jhonyy9):

\[2(x+3)(x-7)=2(x ^{2}-4x-21)=2x ^{2}-8x-42\]